OpenStudy (kj4uts):

Determine the average rate of change of the function f(x)=-2x^2+5 over the interval from x=1 to x=4. My choices are: -10 10 -8 8 None of the others

OpenStudy (kj4uts):

Do I plug this in like this: -2(1)^2+5=3 -2(4)^2+5=-27

OpenStudy (anonymous):

yes that's a start. You found f(1) = 3 and f(4) = -27. So now you do \[\frac{ -27-3 }{ 4-1 }\]because the equation is\[avg~rate~of~change=\frac{ f(b)-f(a) }{ b-a }\]

OpenStudy (kj4uts):

oh ok I got -10 thank you for explain that with the formula.

OpenStudy (anonymous):

yeah that's right. you're welcome